OpenStudy (anonymous):

how many numbers from 1- 1000 are divisible by both 2 AND 5? Pls help!! @TrashMan914

jimthompson5910 (jim_thompson5910):

Hint: to be "divisible by both 2 AND 5" is to be divisible by 2*5 = 10

jimthompson5910 (jim_thompson5910):

so just list out the multiples of 10 from 1 to 1000 or you can use the idea that 10 = 10*1 20 = 10*2 30 = 10*3 ... .... 980 = 98*10 990 = 99*10 1000 = 100*10 so that will speed up the counting process

OpenStudy (anonymous):

so 1000/10?

jimthompson5910 (jim_thompson5910):

Oh i guess I should stay consistent 10 = 10*1 20 = 10*2 30 = 10*3 ... .... 980 = 10*98 990 = 10*99 1000 = 10*100

jimthompson5910 (jim_thompson5910):

can you see how many numbers there are using that trick?

OpenStudy (anonymous):

100 thanks! btw if you were to ask for the numbers from 1-1000 divisible by both 6 and 9 would you find the least common multiple of 18? the 1000/18?

jimthompson5910 (jim_thompson5910):

yes you would find the LCM of 6 and 9, which is 18, then list out the multiples

OpenStudy (anonymous):

do you have to list out the multiples or can you just do 1000/18=55.5? then get 55 as your answer?

jimthompson5910 (jim_thompson5910):

that works as well since 55*18 = 990 and that's as high as you can go so basically 18 = 18*1 36 = 18*2 ... ... 990 = 18*55

OpenStudy (anonymous):

thank you so much for your help :)

jimthompson5910 (jim_thompson5910):

sure thing